LABORATORY AND ASTRONOMICAL DETECTION OF THE NEGATIVE MOLECULAR ION C_{3}N^{−}

\underline{CARL A. GOTTLIEB}, PATRICK THADDEUS, SANDRA BRÜNKEN, MICHAEL C. Mc-CARTHY, Harvard-Smithsonian Center for Astrophysics; and School of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138; MARCELINO AGÚNDEZ, JOSE CER-NICHARO, Department of Molecular and Infrared Astrophysics, Instituto de Estructura de la Materia, CSIC, 28006 Madrid, Spain; MICHEL GUÉLIN, Insitut de Radioastronomie Millimétrique, 38406 St. Martin d’Hères; and LERMA/Ecole Normale Superieure, 75231 Paris, France.

The negative molecular ion C_{3}N^{−} has been detected at millimeter wavelengths in a low pressure laboratory discharge, and then with frequencies in hand in the molecular envelope of IRC+10216. Spectroscopic constants derived from laboratory measurements of 12 transitions between 97 and 378 GHz allow the rotational spectrum to be calculated well into the submillimeter-wave band. Four transitions of C_{3}N^{−} were detected in IRC+10216 with the IRAM 30 m telescope at precisely the frequencies calculated from the laboratory measurements. The column density of C_{3}N^{−} is 0.5% that of C_{3}N, or approximately 20 times greater than C_{4}H^{−} relative to C_{4}H. The C_{3}N^{−} abundance in IRC+10216 is compared with chemical model calculations,^{a} and observations in TMC–1 with the NRAO 100 m Green Bank Telescope (GBT) are discussed. The fairly high concentration of C_{3}N^{−} achieved in the laboratory implies that other molecular anions containing the CN group may be within reach.

aS. Petrie and E. Herbst, Astrophys. J. Lett. 491, 210 (1997); E. Herbst and Y. Osamura, Astrophys. J., in press (2008).